OpenStudy (anonymous):

A cyclist maintains a constant velocity of 4 m/s headed away from point A. At some initial time, the cyclist is 253 m from point A. What will be his displacement from his starting position after 64 s? Answer in units of m What will be his position from point A after that time? Answer in units of m

OpenStudy (amistre64):

whenever you see postion and velocity; you should know that the equations are related. the derivative of position is velocity

OpenStudy (amistre64):

v(t) = 4 p(t) = 4t + C ; at t=0, i believe it says we are -253 away p(0) = 4(0) + C = -253 C = -253 p(t) = 4t -253

OpenStudy (amistre64):

displacement tho is just distance away from starting point; in this case 4(64) = 256 m ... which is 3 m away from A

OpenStudy (amistre64):

the equation stuff is just extra that wasnt needed :)

OpenStudy (anonymous):

Thanks ALOT